OpenStudy (pulsified333):

17. Assume that the car lot contains 55 percent Porches, 25 percent Corvettes, and 20 percent Lincolns. Of the Porches, 70 percent have two airbags, 80 percent of the Corvettes have two airbags, and 60 percent of the Lincolns have two airbags. You are assigned a car at random. If the car has two airbags, what is the probability that it is not a Porche?

OpenStudy (anonymous):

You are only using two airbag cars, so you need to find the amount of cars with airbags in parking lot. Porche=55/100 * 70/100= 3850/100000 or 38.5% Corvette=25/100 * 80/100= 2000/10000 or 20% Lincoln=20/100 * 60/100= 1200/10000 or 12% These add up to 70.5 which you multiply by (100/70.5) Multiply the percents by the (100/70.5) Subtract the Porche answer from the total 100% The remainder is your answer Sorry if there is a more simple way to do it.